Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. It started only two jobs during March—Job P and Job Q. Job P was completed and sold by the end of March and Job Q was incomplete at the end of March. The company uses a plantwide predetermined overhead rate based on direct labor-hours. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March): Estimated total fixed manufacturing overhead $ 15,600 Estimated variable manufacturing overhead per direct labor-hour $ 1.80 Estimated total direct labor-hours to be worked 3,900 Total actual manufacturing overhead costs incurred $ 22,100 Job P Job Q Direct materials $ 15,000 $ 9,900 Direct labor cost $ 54,000 $ 9,000 Actual direct labor-hours worked 3,000 500

1. What is the amount of underapplied or overapplied overhead?

2. Calculate the cost of goods manufactured using the indirect method.
Answer #1

1) Overhead rate = (15600/3900+1.80) = 5.80 per labor hour

Applied overhead = 3500*5.8 = 20300

Actual overhead = 22100

Under applied overhead = 20300-22100 = 1800

2) Cost of goods manufactured

Beginning work in process 0
Direct material 24900
Direct labor 63000
Overhead 20300
Total manufacturing cost 108200
Less: Ending work in process -21800
Cost of goods manufactured 86400